# UniEdit: A Unified Tuning-Free Framework for Video Motion and Appearance Editing

### 24 Feb 2024 v3

@@ 1 Introduction
The advent of pre-trained diffusion-based [22; 46] text-to-image generators [42; 43; 41] has revolutionized the fields of design and filmmaking, opening new vistas for creative expression. These advancements, underpinned by seminal works in text-to-image synthesis, have paved the way for innovative text-guided editing techniques for both images [35; 20; 4; 5] and videos [58; 6; 33; 62; 13; 39]. Such techniques not only enhance creative workflows but also promise to redefine content creation within these industries.
Video editing, in contrast to image editing, introduces the intricate challenge of ensuring frame-wise consistency. Efforts to address this challenge have led to the development of methods that leverage shared features and structures with the source video [6; 33; 31; 62; 39; 7; 26; 55; 14] through an inversion-then-generation pipeline [35; 46], exemplified by Pix2Video’s approach [6] to consistent appearance editing across frames. To transfer the edited appearance from the anchor frame to the remaining frames consistently, it employs a pre-trained image generator and extends the self-attention layers to cross-frame attention to generate each remaining frame. Despite these advancements in performing video appearance editing (e.g., stylization, object appearance replacement, etc.), these methodologies fall short in editing video motion (e.g., replacing the movement of playing guitar with waving), hampered by a lack of motion priors and limited control over inter-frame dependencies, underscoring a critical gap in video editing capabilities.
Previous attempts [58; 37] at video motion editing through fine-tuning a pre-trained generator on the given source video and then editing motion through text guidance. Although effective, they necessitate a delicate balance between the generative prowess of the model and the preservation of the source video’s content. This compromise often leads to restricted motion diversity and unwanted content variations, indicating a pressing need for a more robust solution.
In response, our work aims to explore a tuning-free framework that adeptly navigates the complexities of editing both the motion and appearance of videos. To achieve this, we identify three technical challenges: 1) it is non-trivial to incorporate the text-guided motion into the source content, as directly applying video appearance editing [39; 14] or image editing [5] schemes leads to undesirable results (as shown in Fig. 5); 2) non-edited content preservation of the source video; 3) inheriting the spatial structure of the source video during appearance editing.
@@ 2.1 Video Generation
Researchers have achieved video generation with generative adversarial networks [51; 44; 54], language models [60; 63], or diffusion models [24; 45; 23; 19; 3; 53; 64; 15; 56; 8]. To make the generation more controllable, recent endeavors have also incorporated additional structure guidance (e.g., depth map) [12; 9; 67; 10; 16; 57], or conducted customized generation [58; 59; 28; 66; 52; 34]. These models have generally learned real-world video distribution from large-scale data, and achieved promising results on text-to-video or image-to-video generation. Based on their success, we leverage the learned prior in the pre-trained model to achieve tuning-free video motion and appearance editing.
@@ 2.2 Video Editing
For appearance editing [62; 11; 13; 29], like turn a video into the style of Van Gogh, the main challenge is to achieve temporal-consistent generation across different frames. Early attempts [6; 31; 39; 7; 26; 55] leveraged text-to-image models with inter-frame propagation to ensure consistency. For instance, Pix2Video [6] replaces the key and value of the current frame with those of the first and previous frame. Video-P2P [33] achieved local editing via video-specific fine-tuning and unconditional embedding optimization [36]. Follow-up studies [14; 62; 38] also leveraged the edit-then-propagate framework with neatest-neighbor field [14], estimated optical flow [62], or temporal deformation field [38]. Despite the promising results, due to the constraint on the source video structure, these approaches are specialized in appearance editing and can not be applied to motion editing directly.
Recent studies have also explored video motion editing with text guidance [58; 37], user-provided motion [27; 47; 11], or specific motion representation [48; 30; 18]. For example, Dreamix [37] proposed fine-tuning a pre-trained text-to-video model with mixed video-image reconstruction objectives for each source video. Then the editing is realized by conditioning the fine-tuned model on the given target prompt. MoCA [61] decoupled the video into the first-frame appearance and the optical flow, and trained a diffusion model to generate video conditioned on the first frame and the text. However, it struggled to preserve the non-edited motion (e.g., background dynamics) as it generates the entire motion from the text. Different from the aforementioned approaches that require fine-tuning or user-provided motion input, we are the first to achieve tuning-free motion and appearance editing with text guidance only.
@@ 3 Preliminaries: Video Diffusion Models
Our proposed UniEdit is built upon video diffusion models. Therefore, we first recap the architecture that is used in common text-guided video diffusion models [56; 2].
@@ 3.1 Overall Architecture
During inference, given standard Gaussian distribution $z_{T}\sim\mathcal{N}(0,1)$, the trained denoising UNet is used to perform $T$ denoising steps to obtain the outputs [22; 46]. If the model is trained in latent space [42], a decoder is employed to reconstruct videos from the latent domain.
@@ 4.1 Tuning-Free Video Motion Editing
One of the key challenges of editing tasks is to inherit the original content (e.g., textures and background) in the source video. To this end, we introduce an auxiliary reconstruction branch. The reconstruction path starts from the same inversed latent $z_{T}$ similar to the main editing path, and then conducts the denoising process with the pre-trained UNet conditioned on the source prompt $P_{s}$ to reconstruct the original frames. As verified in image editing [49; 20; 5], the attention features in the denoising model during reconstruction contain the content of the source video. Hence, we inject attention features of the reconstruction path into the main editing path on spatial self-attention (SA-S) layers for content preservation. At denoising step $t$, the attention operation of the $l$-th SA-S module in the main editing path is formulated as:
@@
where $Q$, $K$, $V$ are features in the main editing path, $V^{r}$ refer to the value feature of the corresponding SA-S layer in the reconstruction branch, $t_{0}$ and $l_{0}$ are hyper-parameters following previous work [5]. By replacing the value of spatial features, the video synthesized by the main editing path retains the non-edited characters (e.g., identity and background) of the source video, as exhibited in Fig. 7. Unlike previous video editing works [31; 25] which introduces a cross-frame attention mechanism (i.e., using the key and value of the first/last frame), we implement Eq. 2 frame-wisely to better tackle source video with large dynamics.
@@ 4.2 Tuning-Free Video Appearance Editing
Previous approaches on video appearance editing [62; 14] mainly realize spatial structure control with the assistance of additional network [65]. When the auxiliary control model fails, it may result in inferior performance in preserving the structure of the original video. Alternatively, we suggest extracting the layout information of the source video from the reconstruction branch. Intuitively, the attention maps in spatial self-attention layers encode the structure of the synthesized video, as verified in Fig. 6. Hence, we replace the query and key of SA-S module in the main editing path with those in the reconstruction branch:
@@ 4.3 Mask-Guided Coordination
To further improve background consistency, we suggest leveraging the foreground/background segmentation mask $M$ to guide the denoising process. There are two possible ways to obtain the mask $M$: the attention maps of CA-S modules with a threshold [20]; or employing an off-the-shelf segmentation model [32] on the source and generated videos. The obtained segmentation masks can be leveraged to 1), alleviate the indistinction in foreground and background; 2), improve the consistency of content between edited and source videos.
@@ 4.4 T2V Models are Zero-Shot TI2V Generators
To make our framework more flexible, we further derive a method to incorporate images as input and synthesize high-quality video conditioned on both image and text-prompt. Different from image animation techniques [17], our method allows the user to guide the animation process with text prompts. Concretely, we first achieve text-to-image (I2V) generation by: 1) transforming input images with simulated camera movement to form a pseudo-video clip [37] or 2) leveraging existing image animation approaches (e.g., SVD [2], AnimateDiff [17]) to synthesis a video with random motion (which may not consistent with the text prompt). Then, we perform text-guided editing with UniEdit on the vanilla video to obtain the final output video.

### 07 Apr 2024 v4

@@ 1 Introduction
The advent of pre-trained diffusion-based [25; 49] text-to-image generators [45; 46; 44] has revolutionized the fields of design and filmmaking, opening new vistas for creative expression. These advancements, underpinned by seminal works in text-to-image synthesis, have paved the way for innovative text-guided editing techniques for both images [38; 23; 4; 5] and videos [61; 6; 36; 65; 16; 42]. Such techniques not only enhance creative workflows but also promise to redefine content creation within these industries.
Video editing, in contrast to image editing, introduces the intricate challenge of ensuring frame-wise consistency. Efforts to address this challenge have led to the development of methods that leverage shared features and structures with the source video [6; 36; 34; 65; 42; 7; 29; 58; 17] through an inversion-then-generation pipeline [38; 49], exemplified by Pix2Video’s approach [6] to consistent appearance editing across frames. To transfer the edited appearance from the anchor frame to the remaining frames consistently, it employs a pre-trained image generator and extends the self-attention layers to cross-frame attention to generate each remaining frame. Despite these advancements in performing video appearance editing (e.g., stylization, object appearance replacement, etc.), these methodologies fall short in editing video motion (e.g., replacing the movement of playing guitar with waving), hampered by a lack of motion priors and limited control over inter-frame dependencies, underscoring a critical gap in video editing capabilities.
Previous attempts [61; 40] at video motion editing through fine-tuning a pre-trained generator on the given source video and then editing motion through text guidance. Although effective, they necessitate a delicate balance between the generative prowess of the model and the preservation of the source video’s content. This compromise often leads to restricted motion diversity and unwanted content variations, indicating a pressing need for a more robust solution.
In response, our work aims to explore a tuning-free framework that adeptly navigates the complexities of editing both the motion and appearance of videos. To achieve this, we identify three technical challenges: 1) it is non-trivial to incorporate the text-guided motion into the source content, as directly applying video appearance editing [42; 17] or image editing [5] schemes leads to undesirable results (as shown in Fig. 5); 2) non-edited content preservation of the source video; 3) inheriting the spatial structure of the source video during appearance editing.
@@ 2.1 Video Generation
Researchers have achieved video generation with generative adversarial networks [54; 47; 57], language models [63; 66], or diffusion models [27; 48; 26; 22; 3; 56; 67; 18; 59; 8]. To make the generation more controllable, recent endeavors have also incorporated additional structure guidance (e.g., depth map) [15; 9; 70; 10; 19; 60], or conducted customized generation [61; 62; 31; 69; 55; 37]. These models have generally learned real-world video distribution from large-scale data, and achieved promising results on text-to-video or image-to-video generation. Based on their success, we leverage the learned prior in the pre-trained model to achieve tuning-free video motion and appearance editing.
@@ 2.2 Video Editing
For appearance editing [65; 14; 16; 32; 11], like turn a video into the style of Van Gogh, the main challenge is to achieve temporal-consistent generation across different frames. Early attempts [6; 34; 42; 7; 29; 58] leveraged text-to-image models with inter-frame propagation to ensure consistency. For instance, Pix2Video [6] replaces the key and value of the current frame with those of the first and previous frame. Video-P2P [36] achieved local editing via video-specific fine-tuning and unconditional embedding optimization [39]. Follow-up studies [17; 65; 41] also leveraged the edit-then-propagate framework with neatest-neighbor field [17], estimated optical flow [65], or temporal deformation field [41]. Despite the promising results, due to the constraint on the source video structure, these approaches are specialized in appearance editing and can not be applied to motion editing directly.
Recent studies have also explored video motion editing with text guidance [61; 40], user-provided motion [30; 50; 14], or specific motion representation [51; 33; 21]. For example, Dreamix [40] proposed fine-tuning a pre-trained text-to-video model with mixed video-image reconstruction objectives for each source video. Then the editing is realized by conditioning the fine-tuned model on the given target prompt. MoCA [64] decoupled the video into the first-frame appearance and the optical flow, and trained a diffusion model to generate video conditioned on the first frame and the text. However, it struggled to preserve the non-edited motion (e.g., background dynamics) as it generates the entire motion from the text. Different from the aforementioned approaches that require fine-tuning or user-provided motion input, we are the first to achieve tuning-free motion and appearance editing with text guidance only.
@@ 3 Preliminaries: Video Diffusion Models
Our proposed UniEdit is built upon video diffusion models. Therefore, we first recap the architecture that is used in common text-guided video diffusion models [59; 2].
@@ 3.1 Overall Architecture
During inference, given standard Gaussian distribution $z_{T}\sim\mathcal{N}(0,1)$, the trained denoising UNet is used to perform $T$ denoising steps to obtain the outputs [25; 49]. If the model is trained in latent space [45], a decoder is employed to reconstruct videos from the latent domain.
@@ 4.1 Tuning-Free Video Motion Editing
One of the key challenges of editing tasks is to inherit the original content (e.g., textures and background) in the source video. To this end, we introduce an auxiliary reconstruction branch. The reconstruction path starts from the same inversed latent $z_{T}$ similar to the main editing path, and then conducts the denoising process with the pre-trained UNet conditioned on the source prompt $P_{s}$ to reconstruct the original frames. As verified in image editing [52; 23; 5], the attention features in the denoising model during reconstruction contain the content of the source video. Hence, we inject attention features of the reconstruction path into the main editing path on spatial self-attention (SA-S) layers for content preservation. At denoising step $t$, the attention operation of the $l$-th SA-S module in the main editing path is formulated as:
@@
where $Q$, $K$, $V$ are features in the main editing path, $V^{r}$ refer to the value feature of the corresponding SA-S layer in the reconstruction branch, $t_{0}$ and $l_{0}$ are hyper-parameters following previous work [5]. By replacing the value of spatial features, the video synthesized by the main editing path retains the non-edited characters (e.g., identity and background) of the source video, as exhibited in Fig. 7. Unlike previous video editing works [34; 28] which introduces a cross-frame attention mechanism (i.e., using the key and value of the first/last frame), we implement Eq. 2 frame-wisely to better tackle source video with large dynamics.
@@ 4.2 Tuning-Free Video Appearance Editing
Previous approaches on video appearance editing [65; 17] mainly realize spatial structure control with the assistance of additional network [68]. When the auxiliary control model fails, it may result in inferior performance in preserving the structure of the original video. Alternatively, we suggest extracting the layout information of the source video from the reconstruction branch. Intuitively, the attention maps in spatial self-attention layers encode the structure of the synthesized video, as verified in Fig. 6. Hence, we replace the query and key of SA-S module in the main editing path with those in the reconstruction branch:
@@ 4.3 Mask-Guided Coordination
To further improve background consistency, we suggest leveraging the foreground/background segmentation mask $M$ to guide the denoising process [13; 12]. There are two possible ways to obtain the mask $M$: the attention maps of CA-S modules with a threshold [23]; or employing an off-the-shelf segmentation model [35] on the source and generated videos. The obtained segmentation masks can be leveraged to 1), alleviate the indistinction in foreground and background; 2), improve the consistency of content between edited and source videos.
@@ 4.4 T2V Models are Zero-Shot TI2V Generators
To make our framework more flexible, we further derive a method to incorporate images as input and synthesize high-quality video conditioned on both image and text-prompt. Different from some image animation techniques [2], our method allows the user to guide the animation process with text prompts. Concretely, we first achieve image-to-video (I2V) generation by: 1) transforming input images with simulated camera movement to form a pseudo-video clip [40] or 2) leveraging existing image animation approaches (e.g., SVD [2], AnimateDiff [20]) to synthesis a video with random motion (which may not consistent with the text prompt). Then, we perform text-guided editing with UniEdit on the vanilla video to obtain the final output video.